### v3

# FLUKE: A Linguistically-Driven and Task-Agnostic Framework for Robustness Evaluation

Abstract
We present FLUKE (framework for linguistically-driven and task-agnostic robustness evaluation), a framework for assessing model robustness through systematic minimal variations of test data. FLUKE introduces controlled variations across linguistic levels — from orthography to dialect and style — and leverages large language models (LLMs) with human validation to generate modifications. We demonstrate FLUKE’s utility by evaluating both fine-tuned models and LLMs across six diverse NLP tasks (four classification and two generation tasks), and reveal that (1) the impact of linguistic variations is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) LLMs still exhibit significant brittleness to certain linguistic variations, with reasoning LLMs surprisingly showing less robustness on some tasks compared to base models, and scaling improving robustness only for surface-level modifications; (3) models are overall more brittle to natural, fluent modifications such as syntax or style changes (and especially to negation), compared to corruption-style tests such as letter flipping; (4) the ability of a model to use a linguistic feature in generation does not correlate to its robustness to this feature on downstream tasks. These findings highlight the importance of systematic robustness testing for understanding model behaviors.
FLUKE: A Linguistically-Driven and Task-Agnostic Framework for Robustness Evaluation
Yulia Otmakhova1∗    Hung Thinh Truong1∗    Rahmad Mahendra 3 Zenan Zhai 2 Rongxin Zhu 2    Daniel Beck 4    Jey Han Lau1 1The University of Melbourne    2Oracle    3Universitas Indonesia    4RMIT University {y.otmakhova, thinh.truong}@unimelb.edu.au    rahmad.mahendra@cs.ui.ac.id zenanzhai@gmail.com    rongxin.zhu@unimelb.edu.au, daniel.beck@rmit.edu.au,     jeyhan.lau@unimelb.edu.au
1 Introduction
The standard evaluation paradigm for machine learning models is to assess their performance on test data that shares the same distribution to the training set. Language evaluation benchmarks (Nie et al. (2020); Zellers et al. (2019); Wang et al. (2018, 2019); inter alia) are typically created based on such an assumption [^1] An exception here is benchmarks specifically designed for evaluating large language models, which typically have no training data (Zhong et al., 2024; Hendrycks et al., 2021; Srivastava et al., 2022). and over the years we have seen a trend of increasing model performance; in some instances the state-of-the-art models are on par with human performance (Hupkes et al., 2023; Le Bras et al., 2020). Despite this positive trend, many studies found that models still perform poorly when dealing with minor variations (Ribeiro et al., 2020; McCoy et al., 2019), minimally contrastive pairs (Sennrich, 2017; Warstadt et al., 2020), out-of-domain data (Singhal et al., 2023), and adversarial input (Jia and Liang, 2017; Ebrahimi et al., 2018). This reveals that the success of NLP models may not be solely attributed to improved linguistic understanding and that they may have learned the data rather than the task (Li et al., 2023; Hupkes et al., 2023; Linzen, 2020).
In this paper, we introduce FLUKE (framework for linguistically-driven and task-agnostic robustness evaluation), a task-agnostic framework to generate minimal modifications of existing test data based on linguistic variations to evaluate model robustness. [^2] Under the generalisation taxonomy of Hupkes et al. (2023), our modifications target specifically covariate shifts. That is, an original input (everyone loves dogs) is modified to examine a model’s robustness in handling linguistic variations such as orthography (evryone loves dogs), syntax (dogs are loved by everyone), semantics (no one loves dogs), etc. These modifications are generated by prompting a large language model (LLM), and then validated manually to assess whether they target the intended linguistic property and require updating the ground-truth labels.
To demonstrate the generalisability of FLUKE, we apply it to modify instances for four classification tasks of various complexity — coreference resolution, dialogue contradiction detection, named entity recognition (NER), and sentiment analysis — as well as two generation tasks — instruction-following evaluation (IFEval) and mathematical reasoning (GSM). We then test a number of models — including fine-tuned pre-trained models (PLMs: BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019) and T5 (Raffel et al., 2020)), base LLMs (GPT-4o, Llama 3.1-405b, and Claude 3.5), as well as reasoning LLMs (GPT-5 and DeepSeek R1) — and examine their predictions in the original and modified instances. When measuring the overall robustness of the model, we account for two cases: where the label of the instance is supposed to stay the same after modification and where the label changes. Accordingly, we consider the model to be robust if its predictions either remain the same or change together with the label. We find that: (1) Robustness to modifications highly depends on the task. (2) Even though LLMs are more robust than PLMs, they are still brittle to many modifications, and for some tasks the brittleness is surprisingly more prominent in reasoning LLMs. (3) Scaling helps to improve robustness only to surface-level modifications. (4) Natural, linguistically equivalent modifications are more damaging than traditional adversarial modifications such as random punctuation insertion; and (5) Models can be robust to linguistic features that they struggled to understand when generating, and brittle to features that are easy for them to employ. Thus, using only a handful of well-known tests may fail to reveal the brittleness of a model, highlighting the necessity of a comprehensive, task-agnostic testing framework like FLUKE. To summarise, our contributions are:
  • We present FLUKE, a task-agnostic framework to generate minimal modifications of test data based on linguistic variations. We design prompts to automate the modifications, and perform extensive quality checks, showing that this approach works well without substantial human effort for most modifications.
  • We demonstrate the utility of FLUKE by evaluating the performance of multiple NLP models, and show that a comprehensive, linguistically driven approach helps to reveal their task-specific shortcomings.
  • We release our prompt templates, evaluation scripts, and validated modifications, to enable other researchers to test their models and tasks. [^3] https://fluke-nlp.github.io/ We also publish a dashboard that allows to assess models performance by comparing their outputs on original and modified instances. [^4] https://fluke-viewer.pages.dev/
2 Related Work
Task-based Benchmarking
Much of previous work in evaluating models is centered around collating datasets that are representative of a range of typical NLP applications (Wang et al., 2018; Song et al., 2023). The datasets in these benchmarks usually follow the typical paradigm of having an in-domain test set, coming from the same distribution as the training data. While this is useful in many experimental settings (such as model comparison), it is restricted to reporting in-domain performance. The importance of evaluation in out-of-domain scenarios is well known. Many shared tasks in NLP include test sets that are unrelated to the training data (Pei et al., 2023; Ousidhoum et al., 2024, inter alia). The development of so-called “challenge sets”, initially explored in machine translation (Goyal et al., 2022) but now extended to LLMs (Zhong et al., 2024) also falls into this space. However, the range of domain differences is massive, which explains why it is hard to create “general”, collated benchmarks for out-of-domain scenarios.
Linguistic Competence Probing
Another way to assess models is through linguistic competence. A common approach is to use probing classifiers, tuned on datasets tailored to specific competences, such as morphology or syntax (Blevins et al., 2023; Mahowald et al., 2024). Holmes (Waldis et al., 2024) attempt to unify multiple datasets created to test linguistic competences. While probing for linguistic competence can provide insights about the model’s understanding of specific linguistic phenomena, it is important to note that model competence does not necessarily translate to task performance, as we show below in Section˜6.5.
Robustness Evaluation
One line of work explores the idea of minimal contrastive pairs (Kann et al., 2019; Marvin and Linzen, 2018; Warstadt et al., 2019; Linzen et al., 2016), to evaluate a model’s robustness to a specific phenomena in the context of a specific task (e.g. subject-verb agreement in syntax (Warstadt et al., 2020) or novel words in machine translation (Sennrich, 2017)). Another line focuses on creating adversarial examples to understand model failures (Eger and Benz, 2020; Keller et al., 2021; Tan et al., 2020; Formento et al., 2023; Li et al., 2024). These studies use model loss or gradient to learn how to create perturbations of test data that break the model.
Ribeiro et al. (2020) introduce the Checklist framework to test linguistic capabilities, but it is not fully task-agnostic, as it requires developers to design their own suite of tests (which, as we show below, may not fully reveal model brittleness), and uses templates to create modifications, which limits their diversity. To address these limitations, we propose a comprehensive suite of varied and natural linguistic modifications together with LLMs prompts that are ready to be applied to any task.
3 FLUKE
FLUKE systematically covers phenomena related to different levels of language structure, as well as some model biases. Simplified examples of each test are in Table 1.
Test Subtest Original Modified Orthography Spelling Addition beautiful beautifull Omission fantastic fantstic Swapping not a bid deal not a big dael Capitalization Upper to lower Battlefield 3 battlefield 3 Lower to upper Hinckley did not hit Reagan Hinckley did not Hit Reagan sPoNgEcAsE The professor The pRoFeSsOr ALL CAPS Michael Phelps MICHAEL Phelps Punctuation Add not exactly the bees knees not exactly, the bees knees Change It’s worth tracking down. It’s worth tracking down! Remove space so little movie so littlemovie Morphology Derivation Derived killed assassinated Compound Compound new brand-new Syntax Voice Active to passive Billy beat Tommy Tommy was beaten by Billy Gramm. role Entity swap Bob sued Bill Bill sued Bob Conjunctions Coordinating excessive hunting excessive hunting and poaching Semantics Concept Synonym suspect doubt Hyper/hyponym organization association Nonce word The bowl had a crack The bowl had a vibble Idiom they were tasty they were a real treat Negation Verbal They were afraid of the robots They were not afraid of the robots Absolute They were afraid of the robots They were afraid of none of the robots Approximate They were afraid of the robots They were seldom afraid of the robots Lexical They were afraid of the robots They were fearless of the robots Double They were afraid of the robots They were not unafraid of the robots Discourse Discourse markers Addition Toyota has Lexus: they are built for the rich. Toyota has Lexus, and they are built for the rich. Change The boss fired the worker when he stopped performing well. The boss fired the worker after he stopped performing well. Remove Tony helped Jeff as he needed help. Tony helped Jeff, he needed help. Appraisal Addition She turns her down. She coldly turns her down. Varieties Style Casual There is no pleasure in watching a child suffer It’s no fun seeing a kid suffer Dialect Singlish He would not say no. He dun wan say no. Biases Temporal Old-fashioned He treats her badly. He treats her ill. Geographical Names Anna tried again Dongxin tried again Cultural entities The bat hit the ball The lakau hit the polo Length Shorten The lion saw the fish and it was swimming The lion saw the fish swimming. Lengthen Joseph did not defeat William Joseph did not manage to defeat William
Table 1: FLUKE: our proposed linguistically-driven tests.
3.1 Linguistic tests
We design tests for all levels of language structure, starting from low-level orthography tests and ending with dialect and style modifications.
Orthography
Here we test the ability of models to deal with changes in spelling (adding or omitting a letter, as well as swapping two letters), capitalization (modifying the register to lower case, Upper case, sPoNgEcAsE or ALL CAPS) and punctuation (where we add or change a punctuation mark, or remove the space between two words).
Morphology
We test if more complex words are harder for models. In derivation modification, we change a non-derived word (a word without any suffixes or prefixes) into a derived word. For compound modifications, we replace a word which has one root with a word that has several stems.
Syntax
Here we test if a model adjusts to changing the sentence structure. In particular, we change a verb from active to passive (the subject and object are flipped accordingly to preserve the meaning). For grammatical role tests, we leave the verb intact, but swap two entities. For conjunctions modification we add an extra word combined with a coordinating conjunction and or or.
Semantics
We devise tests that involve different degrees of meaning change. First, in concept modifications, we replace words with synonyms, with hypernyms or hyponyms (words that have a broader or narrower meaning), with artificially created, non-existing nonce words [^5] We use a list of nonce words from Cremers (2022)., and with idioms. We also use negation to test for the change of meaning that goes beyond individual words. Apart from verbal negation (adding not to the verb), we consider less straightforward types of negation: approximate, where the meaning is only partially negated, absolute, which explicitly excludes non-negative interpretation, as well as lexical negation (affixal negation and antonymy) (Pullum and Huddleston, 2002; Otmakhova et al., 2022). Finally, we look at double negation, i.e. grammatically valid structures which contain two negation elements (Truong et al., 2022).
Discourse
We look at sensitivity to manipulating discourse markers: we add a discourse marker, explicitly connecting ideas, change a discourse marker, which can lead to a different meaning, or delete it, making the text more difficult to interpret. We also test robustness to adding markers of appraisal: sentiment-bearing words or phrases.
Language Varieties
In these tests the changes are more extensive and span across all linguistic aspects. Specifically, we change the style to highly casual English, and rewrite the sentence in a particular dialect: Singaporean creole English (Singlish), which has systematic differences from American or British English varieties used in benchmarks.
3.2 Bias tests
In addition to linguistic tests, we include tests for data artifacts that models are known to exploit. We test if changing words to their old-fashioned variants affects the performance (temporal bias). When testing for geographical bias (Faisal and Anastasopoulos, 2023; Godey et al., 2024), we randomly select from a list of regions that are likely to be underrepresented in training corpora. [^6] Africa, Middle East, Southeast, East and Central Asia, Oceania, Latin America, Eastern Europe Then we change proper nouns to names and locations specific to such regions, and replace common nouns with relevant cultural entities. Finally, in length bias tests we either shorten or lengthen the text.
4 Task Evaluation
We apply modifications to test model robustness for six tasks, including four classification tasks (coreference resolution, dialogue contradiction detection, named entity recognition, sentiment analysis) and two generation tasks (instruction-following evaluation, IFEval, and mathematical reasoning, GSM).
4.1 Test set creation and validation
We generate modifications (Section˜3) for each task, randomly selecting samples from a test partition of the original dataset and prompting GPT-4o with instructions. We create and evaluate a general prompt template for each modification (see Section˜A.1 for an example), and then use the same prompt template for each task with minimal modifications (such as specifying the text type).
We follow several steps to ensure the high quality of modifications. To improve their diversity, we stratify modifications across subtests (for example, different types of negation) and, where possible, randomly select a variable from a list (such as region and country for geographical bias or non-existing word for nonce semantic test) and pass it to the prompt. To ensure that modifications were applied minimally, i.e. the difference in performance can be attributed to a particular linguistic feature change rather than to a large number of changes, we calculate Levenshtein distance between original and modified samples, discarding those which changed more than expected. Finally, we employ thorough human checks to ensure that modifications are valid and meaningful. For classification tasks we ran this validation in two stages on Prolific [^7] https://www.prolific.com/, with thorough quality control and four annotators per sample (see Section˜A.2 for annotation details, inter-annotator agreement and quality statistics). During the first stage, we check if resulting texts have the required modification and are still fluent, rejecting samples that do not meet these criteria. We find that most tests have a high retain rate of 70-90%. During the second stage, we ask annotators to select the tasks’ label (such as positive or negative for sentiment analysis) for modified samples, without knowing the labels for original samples, to check if the ground-truth label changed (e.g. after applying negation the sentiment can change from positive to negative). To make sure the task is solvable after modification, i.e. a label is not ambiguous or hard to choose, we add an option such as “Neutral” for sentiment analysis or “Not sure” for other tasks, and reject samples where annotators chose it.
For generation tasks, to demonstrate how the framework can be applied with minimal human intervention, we skip the first stage of human checks and only use tests with reliable modifications (over 70% retain rate). For the second stage, the authors of the paper validate modified samples in terms of whether the modification changes the label (for example, if it changes the premises in the math task, leading to a different answer) and whether the task is still solvable (for example, if there are modifications that prevent correct instruction following). Examples of such valid and rejected modifications, as well as modifications that led to label change, are shown in Table˜9 in Appendix. Such setup reduces the human effort, making the framework generalizable to new tasks, while still ensuring the validity of modifications.
In total, we produce around 100 modified instances for each test in each task (1700 modified instances for each classification task and 1300 instances per generation task).
4.2 Tested models and evaluation
We experiment with three smaller PLMs (BERT, GPT-2, and T5), three base LLMs (GPT-4o, Llama 3.1 and Claude 3.5), and two reasoning LLMs (GPT-5, DeepSeek R1). Both PLMs and LLMs are tested on classification tasks, while only LLMs are used for generation tasks. The pretrained models are fine-tuned on the respective training set with an additional task-specific classification layer on top (Section˜A.5 in Appendix). The LLMs are prompted with task-related instructions in a zero-shot way (Table˜10).
To ensure a fair comparison of performance between the original and modified set, for each test we evaluate only the subset of the original samples that corresponds to the modified instances. We measure instability via an “unrobustness” score $U$ computed on paired original–modified examples. Let $o_{i},m_{i}$ be per-sample correctness indicators for the original and modified items, respectively. [^8] Correctness is defined as whether the prediction matches the ground truth label. Note that we use this formulation because it handles the situations where there is a label change after modification (e.g. sentiment has flipped after negation is introduced). In those cases, the prediction will also need to change for it to be “correct”. Then
$U=\frac{1}{N}\sum_{i=1}^{N}\big|m_{i}-o_{i}\big|\cdot 100$ (1)
where $N$ is the number of instances. Note that $U$ counts both degradations (1→0) and improvements (0→1) as “unrobust”: our rationale is that if after modification the model correctly predicts a previously incorrect sample, it is still brittle since the model relies on a particular way the text is written rather than understanding its content.
We follow the original metrics used for the task to define correctness. For dialogue contradiction detection, coreference resolution, sentiment analysis, and GSM the correctness metric is accuracy. For NER, we use per-sample entity-level mean F1 (Nakayama, 2018), while IFEval employs strict success (all constraints satisfied) as correctness.
5 Results
5.1 Use Case: NER and Coreference tasks
In this section we present two classification tasks: named entity extraction (NER) and coreference resolution (Coreference). [^9] The results for the other two classification tasks – Sentiment analysis and Dialogue Contradiction resolution – are in Section A.8 due to space constraints. For Coreference task, we use KnowRef dataset (Emami et al., 2019), which consists of pronoun disambiguation problems where each sample has the Winograd Schema Challenge (Levesque et al., 2012) format, consisting of a sentence, the pronoun we want to disambiguate, two candidates, and the index of the correct candidate as label. For NER, we chose Few-NERD (Ding et al., 2021), a dataset of Wikipedia sentences, using its 8 first-level entity categories, i.e., Person, Location, Organization, Art, Building, Product, Event, Misc. As explained in Section˜4.1, we thoroughly check and update the labels to make sure the task is still valid and solvable after modification (for example, check if the pronoun can still be resolved or if the entity has changed).
Interestingly, unrobustness results for NER (see Table˜2) and Coreference (Table˜3) show different, almost orthogonal trends in terms of the tests the models were brittle to. For NER the most revealing test is Geographical bias, where all models predict entities from some locales better than from the others. While in some cases the models struggled to identify entities after modification, in the majority of cases it was actually easier for the models to predict “exotic” entities from non-English locales. Moreover, almost all models (except for BERT) lacked robustness to Punctuation modifications, heavily relying on capitalization as a cue for named entity presence or type. Interestingly, while base LLMs overall are more robust to modifications on this task, reasoning LLMs (especially DeepSeek) show major fluctuations across all NER tests, demonstrating similar unrobustness to PLMs.
Conversely, for the Coreference task, while models were relatively robust to the tests affecting NER, they demonstrated major instability on tests involving modifications of syntactical structure and meaning, such as change of active Voice to passive, swapping Grammatical roles, or introducing Negation, as well as perturbations involving significant changes of context such as Length bias, change to casual Style, or to Singlish dialect. Unlike NER, on Coreference tasks LLMs (with the exception of GPT-4o) perform better than PLMs; however, they still demonstrate unrobustness to the majority of tests, with even the best model (GPT-5) significantly affected by grammatical roles change, negation and style modifications.
Category Modification PLM LLM Avg
BERT GPT2 T5 GPT4o Claude Llama GPT5 DS
Bias Temporal 3.7 3.0 1.2 6.2 4.3 1.8 7.7 10.6 4.8
Geographical 25.1 27.6 29.0 22.3 26.0 27.0 22.0 27.8 25.8
Length 11.8 12.2 13.1 12.7 15.7 6.4 9.7 12.6 11.8
Orthogr. Spelling 4.3 2.5 1.4 6.6 3.0 3.9 7.3 11.4 5.0
Capitalization 0.9 19.3 13.1 11.3 8.9 15.3 9.1 13.6 11.5
Punctuation 6.5 3.7 4.9 7.1 9.1 7.9 7.8 15.0 7.8
Morphol. Derivation 1.9 4.2 5.8 3.7 2.0 1.3 9.5 8.2 4.6
Compound 3.1 0.6 5.2 3.1 1.7 1.0 7.0 12.9 4.3
Syntax Voice 7.8 10.8 5.7 7.5 5.5 4.5 8.3 11.3 7.7
Grammar 8.0 15.9 10.5 8.3 3.5 5.3 10.8 12.4 9.3
Conjunction 9.1 7.6 7.7 9.4 7.5 8.0 11.7 13.2 9.3
Semantics Concept 5.0 8.9 5.3 6.5 4.8 3.9 8.5 9.4 6.5
Negation 4.8 5.2 6.5 8.2 3.9 4.6 10.5 15.2 7.4
Discourse Disc. markers 4.7 1.6 5.4 2.2 1.9 3.2 6.3 8.7 4.2
Appraisal 5.5 2.6 4.8 3.2 3.1 2.9 7.1 14.0 5.4
Varieties Style 11.8 12.2 13.1 3.6 4.8 6.4 7.4 10.9 8.8
Dialectal 12.6 7.4 8.4 5.9 8.9 4.7 7.0 13.2 8.5
Average 7.4 8.6 8.3 7.5 6.7 6.4 9.3 13.0 8.4
Table 2: NER: Unrobustness (U, %) by model and modification. Claude stands for Claude 3.5, Llama for Llama 3.1, and DS for DeepSeek R1. Full results with confidence interval (CI) in Table˜13.
Category Modification PLM LLM Avg
BERT GPT2 T5 GPT4o Claude Llama GPT5 DS
Bias Temporal 9.0 4.0 6.0 8.0 1.0 7.0 3.0 8.0 5.8
Geographical 8.0 14.0 9.0 10.0 4.0 5.0 6.0 10.0 8.2
Length 19.2 15.2 14.1 15.2 20.2 17.2 8.1 13.1 15.3
Orthogr. Spelling 11.2 3.1 7.1 5.1 2.0 4.1 7.1 4.1 5.5
Capitalization 15.2 14.1 7.1 7.1 4.0 5.1 2.0 6.1 7.6
Punctuation 1.0 7.1 1.0 1.0 3.0 1.0 2.0 4.0 2.5
Morphol. Derivation 4.1 3.1 5.1 4.1 1.0 1.0 5.1 2.0 3.2
Compound 6.2 4.2 3.1 5.2 6.2 6.2 3.1 4.2 4.8
Syntax Voice 35.8 34.7 41.1 26.3 15.8 30.5 9.5 15.8 26.2
Grammar 30.6 27.8 19.4 22.2 18.1 19.4 16.7 20.8 21.9
Conjunction 5.2 10.3 8.2 8.2 4.1 7.2 5.2 6.2 6.8
Semantics Concept 8.0 3.0 5.0 11.0 18.0 9.0 10.0 10.0 9.2
Negation 25.5 23.5 24.5 21.4 24.5 22.4 22.4 30.6 24.4
Discourse Disc. markers 8.0 6.0 8.0 19.0 10.0 7.0 9.0 11.0 9.8
Appraisal 5.0 8.0 4.0 8.0 7.0 9.0 5.0 7.0 6.6
Varieties Style 16.0 19.0 18.0 19.0 14.0 14.0 14.0 12.0 15.8
Dialect 8.3 26.9 22.2 24.5 11.8 16.7 2.9 12.7 15.8
Average 12.7 13.2 11.9 12.7 9.7 10.7 7.7 10.5 11.1
Table 3: Coreference: Unrobustness (U, %) by model and modification. Notation as in Table˜2. Full results with CI in Table˜14.
5.2 Use case: GSM and IFEval
We choose two generation tasks — mathematical reasoning and instruction following — to demonstrate the applicability of FLUKE to testing generative models. For the mathematical reasoning task we choose Grade School Math (GSM8K) dataset (Cobbe et al., 2021), which contains mathematical problems expressed in natural language. To simplify the task, we use only the final computed answer (rather than the reasoning steps) to evaluate the solution correctness. For the instruction following task, we use the IFEval benchmark (Zhou et al., 2023) which contains varied generation prompts (“Write me a letter…”) that include 25 types of verifiable instructions such as “include word XXX at least 3 times" or “use 2 bullet points". Due to the nature of these tasks, we evaluate only LLMs on them. We ensure that the tasks are solvable after the modification, and that the expected output (solution or generated text) is still the same for all tests except for some types of Negation (verbal, absolute, lexical) where the modification changes the condition of the task and thus the model is expected to generate a different answer (see Section˜4.1).
We find that on the GSM task (Table˜4) models demonstrate major brittleness to Negation tasks, where all of them fail to take into account the changed or even reversed logic of the premise, showing memorisation effects. Base LLMs such as Claude-3.5 and especially Llama 3.1 are also affected by context changes, such as failing to accomodate for changed names and currencies in Geographical bias test, or explaining the problem in informal style or Singlish dialect. Interestingly, reasoning LLMs are more robust to such changes.
Category Modification GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 1.0 2.0 4.0 0.0 1.0 1.6
Geographical 5.0 5.0 7.0 1.0 2.0 4.0
Length 4.0 2.0 2.0 0.0 2.0 2.0
Orthogr. Spelling 1.0 0.0 2.0 1.0 0.0 0.8
Capitalization 3.0 1.0 5.0 1.0 1.0 2.2
Punctuation 0.0 0.0 5.0 1.0 1.0 1.4
Semantics Concept 1.0 0.0 5.0 2.0 1.0 1.8
Negation 15.0 15.0 18.0 7.0 9.0 12.8
Discourse Appraisal 0.0 0.0 1.0 1.0 1.0 0.6
Varieties Style 4.0 3.0 6.0 4.0 3.0 4.0
Dialect 2.0 4.0 6.0 1.0 2.0 3.0
Syntactic Conjunction 0.0 1.0 4.0 1.0 1.0 1.4
Voice 2.0 2.0 6.0 1.0 2.0 2.6
Average 2.9 2.7 5.5 1.6 2.0 2.9
Table 4: GSM: Unrobustness (U, %) by model and modification. Notation as in table˜2. Full results with CI in Table˜17.
Compared to solving mathematical problems, the LLM’s ability to follow instructions are much more affected by the way the instructions are written — with both base and reasoning LLMs showing significant fluctuations in performance across all tests. This includes not only the “usual suspects” such as Negation, Style and Dialect, but also more innocuous changes such as adding an appraisal word or phrase (please repeat -> please be so nice as to repeat), adding a word with a conjunction (Rewrite the following text -> Rewrite and transform the following text), changing the case of a single word in instruction, etc. Here, reasoning models do not seem to handle modifications better than the base ones.
Category Modification GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 11.1 4.0 12.1 5.1 12.1 8.9
Geographical 9.0 12.0 14.0 9.0 10.0 10.8
Length 11.0 10.0 18.0 6.0 16.0 12.2
Orthogr. Capitalization 8.1 5.1 10.1 9.1 13.1 9.1
Punctuation 7.1 3.0 13.1 8.1 12.1 8.7
Spelling 3.1 4.1 7.2 6.2 8.2 5.8
Syntax Conjunction 11.0 7.0 6.0 8.0 11.0 8.6
Voice 9.0 9.0 10.0 6.0 15.0 9.8
Semantics Concept 9.1 7.1 11.1 7.1 7.1 8.3
Negation 24.0 22.0 24.0 23.0 23.0 23.2
Discourse Appraisal 7.1 4.1 11.2 7.1 6.1 7.1
Varieties Style 18.0 7.0 10.0 9.0 11.0 11.0
Dialect 13.0 13.0 12.0 11.0 15.0 12.8
Average 10.8 8.3 12.2 8.8 12.3 10.5
Table 5: IFEval: Unrobustness (U, %) by model and modification. Notation in Table˜2. Full results with CI in Table˜18.
6 Discussion
6.1 Model brittleness is task-specific
We call our framework task-agnostic, since the same set of prompts can be used to generate a battery of tests applicable to all tasks. This, however, does not mean that all tests are meaningful for all the tasks. As the results above (and in Section˜A.8) show, the tests that actually “fired", or showed substantial drops or increases, are different for each task. Even test such as negation which is known to be difficult for models (Ravichander et al., 2022; Truong et al., 2022) has almost no effect on the NER task. On the other hand, seemingly “simple” tests (capitalization) that have a negligible effect on most of the tasks led to some dramatic drops on the NER and IFEval tasks. Thus, our framework is task-agnostic because it can be applied to reveal lack of robustness on a new task without having a preconceived (and potentially wrong) notion of what tests are meaningful for it.
6.2 LLMs are not always more robust than PLMs, and reasoning enhanced training does not always help
In classification tasks, where PLMs were finetuned on the original dataset, one would expect larger fluctuations in performance compared to LLMs, as modifications are likely to change the data distribution and break the identical distribution assumption that the fine-tuning process is based on. This holds for two out of four classification tasks we tested — Coreference resolution (Table˜3) and Sentiment analysis (Table˜16 in Section˜A.8) — where PLMs are overall less robust to perturbations. However, on two other classification tasks — NER (Table˜2) and Dialogue contradiction detection (Table˜15 in Section˜A.8) — while base LLMs are overall more robust than PLMs, reasoning LLMs are as brittle as the latter. Moreover, while LLMs are on average (across all models and tests) better (see the “Average” row in Table˜20), they are sometimes markedly brittle to a particular test. For example, Claude-3.5 is substantially more affected by concept modifications than any PLMs on Coreference resolution, or GPT-4o affected by grammatical role modifications on Dialogue contradiction task.
Comparing between base and reasoning LLMs across tasks does not reveal a clear winner either. Reasoning seems to improve overall robustness on tasks such as GSM (Table˜4), Coreference (Table˜3) or Sentiment Analysys (Table˜16); however, on other tasks the reasoning models perform similar to base LLMs (IFEval, Table˜5) or even worse than them (NER, Table˜2, and Dialogue contradiction detection, Table˜15).
We also test if a reasoning model (GPT-5) is able to handle the modifications better if it is aware of them. To do that, we include additional context into the prompt where we alert the model that the sample was modified, and explicitly specify the type of modification (e.g. by telling it that the active voice was changed to passive). We found no consistent gain when we do this (see Appendix Table˜12).
6.3 Scaling helps but with diminishing effects and only for superficial modifications
To test if increasing the parameter size helps with robustness, we additionally compare Llama-3.1 405B (which was used in all experiments) with two smaller versions – Llama-3.1-8B and Llama-3.1-70B – on the GSM task (the results in Table˜6). Overall, we observe a large improvement in robustness when going from 8B $\rightarrow$ 70B but observe diminishing effects when going beyond that. Low-level orthographic and syntactic modifications benefit strongly from increased model size. In contrast, more complex perturbations, especially negation, remain highly challenging even at 405B. This suggests that scaling primarily improves surface-level robustness, while deeper semantic changes remain brittle.
Category Modification Llama-8b Llama-70b Llama-405b Avg
Bias Temporal 8.0 5.0 4.0 5.7
Geographical 8.0 3.0 7.0 6.0
Length 8.0 2.0 2.0 4.0
Orthogr. Spelling 10.0 2.0 2.0 4.7
Capitalization 11.0 2.0 5.0 6.0
Punctuation 5.0 1.0 5.0 3.7
Syntax Conjunction 12.0 2.0 4.0 6.0
Voice 11.0 2.0 6.0 6.3
Semantic Concept 12.0 1.0 5.0 6.0
Negation 22.0 17.0 18.0 19.0
Discourse Appraisal 7.0 5.0 1.0 4.3
Varieties Style 16.0 6.0 6.0 9.3
Dialect 10.0 6.0 6.0 7.3
Average 10.8 4.2 5.5 6.8
Table 6: GSM Scaling: Unrobustness (U, %) by Llama model size and modification. Notation in Table˜2. Full results with CI in Table˜19.
6.4 Models are overall less robust to natural, linguistically valid modifications
Some of the work on robustness focuses on corrupting the text in adversarial way, such as shuffling characters or changing the case (Belinkov and Bisk, 2017; Li et al., 2018; Eger and Benz, 2020), while some includes natural, lingustically equivalent perturbations such as synonym replacement (Ren et al., 2019; Jin et al., 2020). Our framework combines both types of tests and allows to compare their efficiency for a particular task. We find that on average linguistically valid modifications reveal unrobustness more often than corrupting the text (i.e. such tests as Orthography) (see “Average” column in Table˜20). The exception here are the NER task where orthography is directly relevant to performance, and, surprisingly, IFEval, where the model often confused the instructions regarding capitalization with capitalization used in the instruction. Thus, the choice of the test type – adversarial corruptions or natural modifications – should also be motivated by the downstream task.
6.5 The ability to generate a modification does not entail robustness against it
Though it is a common practice to use LLMs to generate test data, this raises a question of circularity if the same model (in our case, GPT-4o) is used to modify instances and then tested against them. To address this concern, we compare the quality of modifications generated by GPT-4o (as evaluated by annotators in terms of retain rate, i.e. percentage of modifications that were found correct and fluent) with the robustness of the same model to these modifications. As can be seen from Figure˜1, as well as low and insignificant Pearson correlation, there is no clear relation between these two aspects: while some modifications (Geographical bias, Dialect) are easy for the model to generate, they lead to significant unrobustness; on the other hand, some modifications can be very hard (Derivation, Compounds), but the model is less brittle to them. Thus, the ability of the model to understand a linguistic feature (i.e. apply it when generating) does not straightforwardly translate to its ability to handle this feature on downstream tasks, and the samples that were generated by a model so as to include a particular feature can be used to test robustness to that feature.
Figure 1: Generation Performance vs Robustness (100-U) of GPT-4o
7 Conclusion
In this paper, we present FLUKE, a task-agnostic, linguistically-driven framework that introduces minimal modifications to existing datasets to create a benchmark for assessing robustness. We automate these modifications by prompting GPT-4o, followed by thorough manual validation to ensure data quality. To showcase the generalizability of this framework, we apply it to create modified instances across six tasks covering both classification and generation, and evaluate the performance of smaller fine-tuned models and large language models. Our findings reveal that: (1) the impact of modifications is highly task-dependent, with some tests being critical for certain tasks but irrelevant for others; (2) while LLMs demonstrate better overall robustness compared to PLMs, they still exhibit significant brittleness, and on some tasks reasoning LLMs are less robust than base LLMs; (3) increasing the model size helps only with more surface-level modifications and has diminishing effects; (4) models are overall less robust to linguistically valid modifications than to adversarial corruption, but this again depends on the task; (5) the ability of a model to implement a linguistic feature during generation does not correlate with its robustness to that feature on downstream tasks. We hope that FLUKE provides an alternative approach for evaluating models, and that model developers would consider integrating these results into model cards (Mitchell et al., 2019) when releasing models.
8 Limitations
Although FLUKE aims to cover a wide range linguistics capabilities, we acknowledge that it is by no means exhaustive: there are still many areas that it can be expanded, and the current suite of capability tests should serve as a foundation for model developers to create more task-specific tests.
FLUKE is currently designed for English. Although many of its capability tests are applicable to other languages, our LLM prompts, human validation experiments and results are limited to English. To adopt FLUKE to another language, we recommend performing a thorough two-stage human evaluation on a sample of tasks to estimate the retain rate of modifications and reveal the tests that are likely to cause label change. For subsequent task, a simplified evalation with human-in-the-loop is possible.
Although we generate the modifications with GPT-4o, the process is ultimately not fully automated, as we still require human assessment to validate the modifications. That said, since most of the automatic modifications were shown to be valid and we only need a small number of instances to understand model capability, when FLUKE is used to test a particular model for a particular task, we believe the validation can be done in-house by the model developers.
Previous works have pointed out potential biases in data generated by LLMs, stemming from the lack of its diversity (Ding et al., 2024). In our work, such biases are most likely to occur in the geographical bias modification. To ensure a diverse representation, we compile a list of underrepresented regions as discussed in Section˜3.2, and explicitly specify a region from this list in the prompt to generate samples across the variety of underrepresented geographical locations.
Acknowledgements
This research is supported by Oracle. Lau was supported by Australian Research Council
Appendix A Appendix
A.1 Data generation prompts
An example of generating compound word modification for coreference resolution task.
Find any non-compound (single-root) word in the text below and change it into a compound word (word with several roots). Don’t make change to the pronoun.
Example: "a sequence of ridiculous shooting scenes" -> "a sequence of ridiculous shoot-’em-up scenes"
Example: dull acting -> lacklustre acting
Text: Joe raced against Steven because he thought it would be easy.
Pronoun: he
Modified text: Joe raced against Steven because he thought it would be a no-brainer.
A.2 Annotation process and instructions
We perform a two-staged validation of modified samples using a crowdsource annotation platform Prolific  [^10] https://www.prolific.com/. To ensure high quality of annotations, we use a high per hour rate of 12 euro, which is above the minimal payment rate in the annotators’ country of residence and well above the usual Prolific rate (overall cost: 7233 euro), choose only annotators from English-speaking countries (except for the Dialect test where we only used annotators from Singapore), and with a minimum success rate of 90%, as well as perform quality checks as described below. The human validation experiments has been approved by the ethics board (application ID: anonymised), and the annotators were informed of how their data will be used and had to express their consent.
During Stage 1, the annotators received instructions that described an intended modification (Voice, Negation, Style etc., see Section˜3) and were asked if the sample they see is modified correctly according to these instructions (see Figure 2 for an example of annotation interface). For quality control, we include four control questions, which were randomly sampled from a different modification. We remove annotators who fail to answer more than one of such questions (i.e. selecting “Correct” twice), or select one option too frequently (i.e. selecting “Incorrect” for all questions). Each of the samples is annotated by four annotators, and receives at least two annotations after the quality checks. We filter our samples which were not deemed to be “Correct” by the majority (over 50%) of annotators. In case of tie, we select the answer of the annotator with better performance on the control questions. The inter-annotator agreement in terms of Krippendorf’s $\alpha$ (Krippendorff, ), averaged majority-class agreement (percentage of samples where the majority of annotators chose the same label), and the retain rate (percentage of samples that pass phase 1 annotation) are shown in Table 7. We choose Krippendorf’s $\alpha$ as an agreement metrics to accommodate for the difference in number of annotations per sample. We note, however, that it is not directly applicable here as it assumes that all annotations are provided by a fixed, ordered set of observers and factors in the consistency of their decisions, i.e. measures the annotators behavior agreement rather than agreement of labels. Thus, in our case the Majority rate agreement reveals more about the patterns of annotation agreement, and achieves substantially high levels.
During Stage 2, the annotators were given instructions for one of the four tasks (Sentiment Analysis, Dialogue Contradiction Detection, Coreference Resolution, Named Entity Recognition) and asked to assign a label specific for that task (for example, judge a modified sample as having positive or negative sentiment for the Sentiment Analysis task, see Figure 3). For quality checks, we used unmodified samples from the training data of the task (i.e. samples for which we knew the label); and apply the same filtering criteria in Stage 1. The final label was also selected based on the majority vote. We compared the resulting labels for modified samples with the labels for the original data to check if any inconsistencies were due to the genuine label change resulting from the modification, rather than to annotators disagreement. The inter-annotator agreement results (measured similarly to Stage 1) are presented in Table 8.
Figure 2: The example annotation page for the use case Sentiment Analysis (SA) for the linguistic capability test derivation in Stage 1.
Modification Krippendorff’s $\alpha$ Majority rate Retain rate Bias Temporal 0.33 0.85 0.80 Geographical 0.35 0.84 0.82 Length 0.30 0.83 0.72 Morphology Derivation 0.14 0.79 0.50 Compound 0.39 0.86 0.61 Syntax Voice 0.30 0.83 0.76 Grammar 0.17 0.80 0.59 Conjunction 0.39 0.88 0.84 Semantics Concept 0.37 0.84 0.75 Negation 0.42 0.90 0.73 Pragmatic Disc. markers 0.22 0.78 0.63 Appraisal 0.46 0.89 0.86 Varieties Style 0.53 0.91 0.93 Dialect 0.41 0.87 0.86
Table 7: Phase 1 annotation quality
Figure 3: The example annotation page for the use case Sentiment Analysis (SA) in Stage 2.
Krippendorff’s $\alpha$ Majority rate Label change rate Sentiment Analysis Semantics: Negation 0.54 0.84 0.53 Dialogue Contradiction Detection Syntax: Grammar 0.34 0.81 0.33 Semantics: Negation 0.39 0.80 0.41 Coreference Resolution Syntax: Grammar 0.66 0.91 0.82 Semantics: Negation 0.57 0.79 0.49 NER Bias: Geography 0.36 0.59 1.00 Syntax: Grammar 0.43 0.67 0.94 Syntax: Conjunction 0.36 0.63 1.00
Table 8: Phase 2 annotation quality
A.3 Examples of annotation decisions for generative tasks (simplified quality assurance scenario)
We provide examples of decisions where we kept the modification together with the original label for the task, changed the label, or rejected the modification, in Table˜9.
Task Test Original Modified Keep? Reason
GSM Dialect Reggie, Lynn, and Paisley ran together. Paisley ran 4 miles. Reggie ran 5 times what Paisley ran and 3 miles farther than Lynn. How many miles did Lynn run? Reggie, Lynn, and Paisley go running kaki. Paisley ran 4 miles. Reggie ran 5 times what Paisley ran and 3 miles more than Lynn. Lynn run how many miles? Despite substantial language change, nothing has changed in the premises of the mathematical problem, so we keep the sample.
GSM Negation Miss Albert’s class is composed of 12 boys and 12 girls. One-third of the girls and one-fourth of the boys are on varsity. How many students are not on varsity? Miss Albert’s class is composed of 12 boys and 12 girls. One-third of the girls and one-fourth of the boys are not on varsity. How many students are not on varsity? $\rightarrow$ The modification is valid (the problem can be solved), but leads to a different answer (label).
GSM Concept 3 trees each had 7 blue birds in them. 2 different trees each had 4 blue birds. 1 final tree had 3 blue birds. How many blue birds were in the trees in total? 3 trees each had 7 blue birds in them. 2 different trees each had 4 blue birds. 1 final tree had 3 sprat birds. How many blue birds were in the trees in total? “Blue” was replaced by a nonce (non-existing) word “sprat”. It is impossible to determine if “sprat birds” are blue or not, so the problem cannot be solved.
IFEval Temp. bias Write exactly 4 paragraphs about tips for installing a car seat for moms. Use 2 new lines to separate paragraphs. Start the 4th paragraph with the word elm. Write exactly 4 paragraphs about tips for installing a car seat for moms. Use 2 new lines to separate paragraphs. Commence the 4th paragraph with the word ’elm’. The word “begin” was replaced by a slightly outdated word “commence”, which is still well understandable in the context. The modification is valid and does not change the label.
IFEval Negation Write me a resume for Matthias Algiers. Use words with all capital letters to highlight key abilities, but make sure that words with all capital letters appear less than 10 times. Write me a resume for Matthias Algiers. Use words with all capital letters to highlight key abilities, but make sure that words with all capital letters appear more than 10 times. $\rightarrow$ “Less” was changed to “more", so if the model generates less than 10 words in all caps, it should be considered wrong.
IFEval Conjunct. …Your entire response should be in English, and should not contain any capital letters. …Your entire response should be in English, and should not contain any capital letters or punctuation. The modification is fluent and applied correctly, but it makes the task more difficult (it is hard to avoid any punctuation), so we reject it.
Table 9: Decision process for keeping (✓), rejecting (✗) modification, or changing the label ($\rightarrow$)
A.4 LLM settings
We report the exact endpoint and sampling parameters for LLMs in Table˜10.
Model GPT4o Claude-3.5 Llama-3.1 GPT-5 Deepseek R1 Endpoint gpt-4o-2024-05-13 claude-3.5-sonnet llama-3.1-405b-instruct gpt-5 deepseek-r1 # Params Undisclosed Undisclosed 405B Undisclosed 671B Temperature 0 0 0 1 (default by reasoning model) 1 (default by reasoning model) Max token 1024 1024 1024 4096 4096
Table 10: LLMs settings
Sentiment analysis prompt
Classify the sentiment of the given text. Answer with 1 for positive, 0 for negative.
Text: it’s a charming and often affecting journey.
Answer: 1
Dialogue contradiction prompt
Does the last utterance contradict the dialogue context? Answer with 1 if contradict, 0 if not contradict.
Dialogue context:
agent 0: well, i’m a big aerosmith fan, but i also like country.
agent 1: aerosmith is not bad but i love country so much
Last utterance: agent 1: i am a nascar fan too!
Answer: 1
Coreference resolution prompt
Which candidate does the pronoun refer to? Answer with either 0 or 1.
Text: The sniper shot the terrorist because he was a bad guy.
Pronoun: he
Candidates: 0: The sniper, 1: the terrorist
Answer: 1
Named entity recognition prompt
Extract named entities from the text. Possible entity types: ART, BUILDING, EVENT, LOCATION, ORGANIZATION, OTHER, PERSON, PRODUCT. Reply with the list of entities in the format ["text": entity_span , "label": entity_label,].
Text: Ronald will travel to Iceland.
Answer: [{"text": "Ronald", "value": "PERSON"}, {"text": "Iceland", "value": "LOCATION"}]
A.5 PLM settings
We report the finetuning details for PLMs in Table˜11.
SA: We add a classification head on top of PLMs and fine-tune as a sequence classification task. Input: text, Output: label (0|1)
Dialogue: We add a classification head on top of PLMs and fine-tune as a sequence classification task. Input: dialog_context[SEP]last_utterance, Output: label (0|1)
Coref: We add a classification head on top of PLMs and fine-tune as a sequence classification task. Input: text[SEP]pronoun[SEP]candidates. Ouput: label (0|1)
NER: We add a classification head on top of PLMs and fine-tune as a sequence labeling task. Input: Text, Output: Label (CoNLL format)
Fine-tuning details are in Table˜11.
Model BERT GPT-2 T5 Checkpoint bert-base-cased gpt2 t5-base Param 110M 124M 220M Optimizer AdamW AdamW AdamW Loss CrossEntropyLoss CrossEntropyLoss CrossEntropyLoss SA Batch size 16 32 32 Epoch 3 5 10 Learning rate 2e-5 2e-5 5e-5 Max length 128 128 128 Dialog Batch size 8 8 8 Training step 10000 10000 10000 Learning rate 2e-5 2e-5 2e-5 Max length 512 512 512 Coref Batch size 16 16 16 Epoch 3 3 3 Learning rate 2e-5 2e-5 2e-5 Max length 128 128 128 NER Batch size 32 32 32 Epoch 5 5 5 Learning rate 2e-5 2e-5 2e-5 Max length 128 128 128
Table 11: PLM fine-tuning details
A.6 Explicitly making model aware of the modification
[…]
Category Modification SA COREF DIALOGUE NER GSM IFEVAL AVG Bias Geographical 1.0 4.0 -1.1 0.3 2.0 0.0 1.0 Length 1.0 0.0 -1.0 -1.9 3.0 2.0 0.5 Temporal -1.0 -2.0 0.0 -2.3 1.0 2.0 -0.4 Orthography Capitalization 0.0 -1.0 -4.2 -3.7 0.0 -2.0 -1.8 Punctuation 0.0 -1.0 -2.0 -0.1 2.0 2.0 0.1 Spelling 1.0 -4.1 3.0 -2.5 1.0 2.0 0.1 Morphology Compound 0.0 0.0 -2.0 -2.7 NA NA -1.2 Derivation 1.1 -2.0 3.2 -5.2 NA NA -0.7 Syntax Conjunction 0.0 -1.0 1.0 -3.9 0.0 2.0 -0.3 Grammar 0.0 -2.8 1.5 -3.7 NA NA -1.2 Voice -1.0 0.0 0.0 0.6 0.0 6.0 0.9 Semantics Concept 0.0 3.0 -5.0 -3.4 -1.0 6.0 -0.1 Negation 0.0 -2.0 3.0 -3.0 4.0 -5.0 -0.5 Discourse Appraisal 0.0 -3.0 0.0 -2.7 0.0 1.0 -0.8 Disc. markers 2.0 3.0 -3.4 1.0 NA NA 0.6 Varieties Dialect -1.0 2.0 -1.9 0.2 1.0 -2.0 -0.3 Style 0.0 -3.0 1.0 0.1 0.0 6.0 0.7 Average Average 0.2 -0.6 -0.5 -1.9 1.0 1.5 -0.2
[…]
A.7 Use Case: NER and Coreference tasks (extended results)
We present the full results of NER and Coreference tasks with confidence interval in Table˜13 and Table˜14.
Category Modification PLM LLM
BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 3.7 [1.4, 6.9] 3.0 [0.4, 6.4] 1.2 [0.3, 2.6] 6.2 [2.5, 10.8] 4.3 [1.1, 8.2] 1.8 [0.3, 4.0] 7.7 [4.4, 11.5] 10.6 [6.4, 15.4] 4.8 [2.1, 8.2]
Geographical 25.1 [19.6, 30.9] 27.6 [21.7, 33.9] 29.0 [23.6, 34.5] 22.3 [17.5, 27.5] 26.0 [19.5, 32.8] 27.0 [21.5, 32.6] 22.0 [16.7, 27.7] 27.8 [22.7, 33.0] 25.8 [20.4, 31.6]
Length 11.8 [6.6, 17.7] 12.2 [7.4, 17.7] 13.1 [8.5, 18.3] 12.7 [7.6, 18.5] 15.7 [9.3, 23.1] 6.4 [3.0, 10.5] 9.7 [5.2, 15.1] 12.6 [7.8, 18.0] 11.8 [6.9, 17.4]
Orthography Spelling 4.3 [1.6, 7.7] 2.5 [0.4, 5.2] 1.4 [0.5, 2.5] 6.6 [3.8, 9.8] 3.0 [0.5, 6.2] 3.9 [1.6, 6.6] 7.3 [4.3, 10.8] 11.4 [7.3, 15.9] 5.0 [2.5, 8.1]
Capitalization 0.9 [0.1, 1.8] 19.3 [13.1, 26.1] 13.1 [8.5, 18.2] 11.3 [6.6, 16.6] 8.9 [4.5, 14.1] 15.3 [9.7, 21.6] 9.1 [5.6, 13.1] 13.6 [9.1, 18.5] 11.5 [7.2, 16.3]
Punctuation 6.5 [3.0, 10.7] 3.7 [1.3, 6.8] 4.9 [2.4, 7.8] 7.1 [3.6, 11.2] 9.1 [5.0, 13.7] 7.9 [4.1, 12.4] 7.8 [4.5, 11.7] 15.0 [10.3, 20.2] 7.8 [4.3, 11.8]
Morphology Derivation 1.9 [0.5, 3.7] 4.2 [1.4, 7.5] 5.8 [2.1, 10.5] 3.7 [1.1, 7.2] 2.0 [0.1, 4.5] 1.3 [0.1, 3.2] 9.5 [4.4, 15.7] 8.2 [4.2, 12.9] 4.6 [1.7, 8.1]
Compound 3.1 [0.8, 6.1] 0.6 [0.0, 1.7] 5.2 [2.0, 9.1] 3.1 [0.8, 6.5] 1.7 [0.0, 4.4] 1.0 [0.0, 2.7] 7.0 [3.4, 11.2] 12.9 [8.2, 18.1] 4.3 [1.9, 7.5]
Syntax Voice 7.8 [3.9, 12.4] 10.8 [6.0, 16.2] 5.7 [3.1, 8.7] 7.5 [3.6, 12.2] 5.5 [2.1, 9.7] 4.5 [1.5, 8.2] 8.3 [4.3, 13.1] 11.3 [7.2, 16.0] 7.7 [4.0, 12.1]
Grammar 8.0 [4.0, 12.9] 15.9 [10.4, 21.8] 10.5 [6.3, 15.5] 8.3 [4.7, 12.5] 3.5 [1.2, 6.6] 5.3 [1.8, 9.6] 10.8 [6.1, 16.1] 12.4 [8.1, 17.0] 9.3 [5.3, 14.0]
Conjunction 9.1 [4.1, 15.4] 7.6 [4.3, 11.3] 7.7 [4.4, 11.4] 9.4 [5.6, 14.0] 7.5 [3.3, 12.9] 8.0 [3.4, 13.7] 11.7 [7.1, 17.1] 13.2 [8.4, 18.7] 9.3 [5.1, 14.3]
Semantics Concept 5.0 [2.3, 8.4] 8.9 [4.3, 14.2] 5.3 [2.6, 8.7] 6.5 [3.2, 10.5] 4.8 [1.2, 9.3] 3.9 [1.2, 7.5] 8.5 [4.1, 13.7] 9.4 [5.7, 13.7] 6.5 [3.1, 10.7]
Negation 4.8 [2.2, 8.0] 5.2 [2.9, 7.9] 6.5 [3.7, 9.9] 8.2 [4.8, 12.2] 3.9 [1.3, 7.0] 4.6 [2.1, 7.6] 10.5 [6.9, 14.7] 15.2 [10.3, 20.6] 7.4 [4.3, 11.0]
Discourse Disc. markers 4.7 [2.0, 8.0] 1.6 [0.1, 3.6] 5.4 [2.8, 8.3] 2.2 [0.8, 3.9] 1.9 [0.4, 4.0] 3.2 [0.8, 6.1] 6.3 [3.4, 9.5] 8.7 [5.2, 12.9] 4.2 [1.9, 7.0]
Appraisal 5.5 [2.9, 8.6] 2.6 [0.8, 5.0] 4.8 [2.4, 7.7] 3.2 [1.4, 5.3] 3.1 [1.0, 5.8] 2.9 [0.9, 5.5] 7.1 [4.2, 10.4] 14.0 [9.2, 19.2] 5.4 [2.9, 8.4]
Varieties Style 11.8 [6.6, 17.7] 12.2 [7.4, 17.7] 13.1 [8.5, 18.3] 3.6 [1.8, 5.7] 4.8 [2.1, 8.0] 6.4 [2.9, 10.6] 7.4 [4.1, 11.2] 10.9 [6.9, 15.3] 8.8 [5.0, 13.1]
Dialectal 12.6 [8.5, 17.3] 7.4 [4.1, 11.4] 8.4 [5.8, 11.4] 5.9 [3.2, 9.3] 8.9 [4.5, 14.0] 4.7 [2.2, 7.9] 7.0 [4.0, 10.6] 13.2 [8.4, 18.3] 8.5 [5.1, 12.5]
Average 7.4 [4.1, 11.4] 8.6 [5.1, 12.6] 8.3 [5.1, 12.0] 7.5 [4.3, 11.4] 6.7 [3.4, 10.8] 6.4 [3.4, 10.0] 9.3 [5.5, 13.7] 13.0 [8.6, 17.9] 8.4 [4.9, 12.5]
Table 13: NER: Unrobustness (U, %) by model and modification
Category Modification PLM LLM Avg
BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 9.0 [4.8, 16.2] 4.0 [1.6, 9.8] 6.0 [2.8, 12.5] 8.0 [4.1, 15.0] 1.0 [0.2, 5.4] 7.0 [3.4, 13.7] 3.0 [1.0, 8.5] 8.0 [4.1, 15.0] 5.8 [2.8, 12.0]
Geographical 8.0 [4.1, 15.0] 14.0 [8.5, 22.1] 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 4.0 [1.6, 9.8] 5.0 [2.2, 11.2] 6.0 [2.8, 12.5] 10.0 [5.5, 17.4] 8.2 [4.4, 15.2]
Length 19.2 [12.6, 28.0] 15.2 [9.4, 23.5] 14.1 [8.6, 22.3] 15.2 [9.4, 23.5] 20.2 [13.5, 29.2] 17.2 [11.0, 25.8] 8.1 [4.2, 15.1] 13.1 [7.8, 21.2] 15.3 [9.6, 23.6]
Orthography Spelling 11.2 [6.4, 19.0] 3.1 [1.0, 8.6] 7.1 [3.5, 14.0] 5.1 [2.2, 11.4] 2.0 [0.6, 7.1] 4.1 [1.6, 10.0] 7.1 [3.5, 14.0] 4.1 [1.6, 10.0] 5.5 [2.5, 11.8]
Capitalization 15.2 [9.4, 23.5] 14.1 [8.6, 22.3] 7.1 [3.5, 13.9] 7.1 [3.5, 13.9] 4.0 [1.6, 9.9] 5.1 [2.2, 11.3] 2.0 [0.6, 7.1] 6.1 [2.8, 12.6] 7.6 [4.0, 14.3]
Punctuation 1.0 [0.2, 5.5] 7.1 [3.5, 13.9] 1.0 [0.2, 5.5] 1.0 [0.2, 5.5] 3.0 [1.0, 8.5] 1.0 [0.2, 5.5] 2.0 [0.6, 7.1] 4.0 [1.6, 9.9] 2.5 [0.9, 7.7]
Morphology Derivation 4.1 [1.6, 10.0] 3.1 [1.0, 8.6] 5.1 [2.2, 11.4] 4.1 [1.6, 10.0] 1.0 [0.2, 5.6] 1.0 [0.2, 5.6] 5.1 [2.2, 11.4] 2.0 [0.6, 7.1] 3.2 [1.2, 8.7]
Compound 6.2 [2.9, 13.0] 4.2 [1.6, 10.2] 3.1 [1.1, 8.8] 5.2 [2.2, 11.6] 6.2 [2.9, 13.0] 6.2 [2.9, 13.0] 3.1 [1.1, 8.8] 4.2 [1.6, 10.2] 4.8 [2.0, 11.1]
Syntax Voice 35.8 [26.9, 45.8] 34.7 [25.9, 44.7] 41.1 [31.7, 51.1] 26.3 [18.5, 36.0] 15.8 [9.8, 24.4] 30.5 [22.2, 40.4] 9.5 [5.1, 17.0] 15.8 [9.8, 24.4] 26.2 [18.7, 35.5]
Grammar 30.6 [21.1, 42.0] 27.8 [18.8, 39.0] 19.4 [12.0, 30.0] 22.2 [14.2, 33.1] 18.1 [10.9, 28.5] 19.4 [12.0, 30.0] 16.7 [9.8, 26.9] 20.8 [13.1, 31.6] 21.9 [14.0, 32.6]
Conjunction 5.2 [2.2, 11.5] 10.3 [5.7, 17.9] 8.2 [4.2, 15.4] 8.2 [4.2, 15.4] 4.1 [1.6, 10.1] 7.2 [3.5, 14.2] 5.2 [2.2, 11.5] 6.2 [2.9, 12.8] 6.8 [3.3, 13.6]
Semantics Concept 8.0 [4.1, 15.0] 3.0 [1.0, 8.5] 5.0 [2.2, 11.2] 11.0 [6.3, 18.6] 18.0 [11.7, 26.7] 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 10.0 [5.5, 17.4] 9.2 [5.1, 16.4]
Negation 25.5 [17.9, 35.0] 23.5 [16.2, 32.8] 24.5 [17.0, 33.9] 21.4 [14.5, 30.5] 24.5 [17.0, 33.9] 22.4 [15.3, 31.7] 22.4 [15.3, 31.7] 30.6 [22.4, 40.3] 24.4 [17.0, 33.7]
Discourse Disc. markers 8.0 [4.1, 15.0] 6.0 [2.8, 12.5] 8.0 [4.1, 15.0] 19.0 [12.5, 27.8] 10.0 [5.5, 17.4] 7.0 [3.4, 13.7] 9.0 [4.8, 16.2] 11.0 [6.3, 18.6] 9.8 [5.4, 17.0]
Appraisal 5.0 [2.2, 11.2] 8.0 [4.1, 15.0] 4.0 [1.6, 9.8] 8.0 [4.1, 15.0] 7.0 [3.4, 13.7] 9.0 [4.8, 16.2] 5.0 [2.2, 11.2] 7.0 [3.4, 13.7] 6.6 [3.2, 13.2]
Varieties Style 16.0 [10.1, 24.4] 19.0 [12.5, 27.8] 18.0 [11.7, 26.7] 19.0 [12.5, 27.8] 14.0 [8.5, 22.1] 14.0 [8.5, 22.1] 14.0 [8.5, 22.1] 12.0 [7.0, 19.8] 15.8 [9.9, 24.1]
Dialect 8.3 [4.4, 15.1] 26.9 [19.4, 35.9] 22.2 [15.4, 30.9] 24.5 [17.2, 33.7] 11.8 [6.9, 19.4] 16.7 [10.7, 25.1] 2.9 [1.0, 8.3] 12.7 [7.6, 20.6] 15.8 [10.3, 23.6]
Average 12.7 [7.9, 20.3] 13.2 [8.3, 20.8] 11.9 [7.4, 19.3] 12.7 [7.8, 20.4] 9.7 [5.7, 16.8] 10.7 [6.4, 18.0] 7.7 [4.1, 14.5] 10.5 [6.1, 17.8] 11.1 [6.7, 18.5]
Table 14: Coreference: Unrobustness (U, %) by model and modification
A.8 Use Case: Dialogue Contradiction Detection and Sentiment Analysis (extended results)
In this section we introduce two additional classification tasks – Dialogue Contradiction Detection and Sentiment Analysis.
We use the DECODE dataset (Nie et al., 2021) to evaluate dialogue understanding capabilities. Each sample is a multi-turn dialogue, where the final turn may be consistent with (0) or contradict (1) the preceding context. For Sentiment Analysis, We chose the SST-2 dataset (Socher et al., 2013), containing around 11K sentences extracted from movies reviews with either 0 (negative) or 1 (positive) sentiment.
As Table˜15 shows, on Dialogue contradiction detection task base LLMs are overall more robust to modifications than PLMs; however, reasoning LLMs on average perform similar to LLMs. All models show major brittleness to negation, and the majority of them also lack robustness to change of grammatical roles, which shows that they tend to rely on surface-level, token consistency between the context and the final turn, rather than on logical coherence. Moreover, LLMs (but to a lesser extent PLMs) demonstrate geographical bias, whuile PLMs are particularly brittle to casual style.
Category Modification BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 8.0 [4.1, 15.0] 6.0 [2.8, 12.5] 7.0 [3.4, 13.7] 2.0 [0.6, 7.0] 5.0 [2.2, 11.2] 1.0 [0.2, 5.4] 3.0 [1.0, 8.5] 2.0 [0.6, 7.0] 4.2 [1.8, 10.0]
Geographical 6.5 [3.0, 13.5] 8.7 [4.5, 16.2] 12.0 [6.8, 20.2] 13.0 [7.6, 21.4] 15.2 [9.3, 23.9] 12.0 [6.8, 20.2] 18.5 [11.9, 27.6] 13.0 [7.6, 21.4] 12.4 [7.2, 20.6]
Length 10.0 [5.5, 17.4] 6.0 [2.8, 12.5] 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 6.0 [2.8, 12.5] 1.0 [0.2, 5.4] 6.0 [2.8, 12.5] 7.0 [3.4, 13.7] 5.6 [2.6, 11.9]
Orthographic Spelling 6.0 [2.8, 12.5] 6.0 [2.8, 12.5] 4.0 [1.6, 9.8] 8.0 [4.1, 15.0] 3.0 [1.0, 8.5] 2.0 [0.6, 7.0] 1.0 [0.2, 5.4] 6.0 [2.8, 12.5] 4.5 [2.0, 10.4]
Capitalization 0.0 [0.0, 0.0] 6.2 [2.9, 13.0] 6.2 [2.9, 13.0] 3.1 [1.1, 8.8] 2.1 [0.6, 7.3] 2.1 [0.6, 7.3] 4.2 [1.6, 10.2] 1.0 [0.2, 5.7] 3.1 [1.2, 8.1]
Punctuation 4.0 [1.6, 9.8] 2.0 [0.6, 7.0] 8.0 [4.1, 15.0] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 1.0 [0.2, 5.4] 5.0 [2.2, 11.2] 2.0 [0.6, 7.0] 3.5 [1.4, 9.0]
Semantic Concept 7.0 [3.4, 13.7] 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 8.0 [4.1, 15.0] 7.0 [3.4, 13.7] 2.0 [0.6, 7.0] 10.0 [5.5, 17.4] 6.0 [2.8, 12.5] 6.1 [2.9, 12.6]
Negation 29.0 [21.0, 38.5] 39.0 [30.0, 48.8] 31.0 [22.8, 40.6] 31.0 [22.8, 40.6] 30.0 [21.9, 39.6] 30.0 [21.9, 39.6] 25.0 [17.5, 34.3] 25.0 [17.5, 34.3] 30.0 [21.9, 39.5]
Pragmatic Appraisal 6.0 [2.8, 12.5] 10.0 [5.5, 17.4] 5.0 [2.2, 11.2] 7.0 [3.4, 13.7] 6.0 [2.8, 12.5] 4.0 [1.6, 9.8] 10.0 [5.5, 17.4] 11.0 [6.3, 18.6] 7.4 [3.8, 14.2]
Genre Style 12.0 [7.0, 19.8] 10.0 [5.5, 17.4] 12.0 [7.0, 19.8] 6.0 [2.8, 12.5] 4.0 [1.6, 9.8] 2.0 [0.6, 7.0] 6.0 [2.8, 12.5] 6.0 [2.8, 12.5] 7.2 [3.7, 13.9]
Dialect 20.8 [14.1, 29.4] 5.7 [2.6, 11.8] 9.4 [5.2, 16.5] 11.3 [6.6, 18.8] 15.1 [9.5, 23.1] 7.5 [3.9, 14.2] 5.7 [2.6, 11.8] 13.2 [8.0, 21.0] 11.1 [6.6, 18.3]
Syntactic Conjunction 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 7.0 [3.4, 13.7] 6.0 [2.8, 12.5] 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 4.0 [1.6, 9.8] 4.0 [1.6, 9.8] 4.1 [1.7, 9.9]
Voice 9.0 [4.8, 16.2] 11.0 [6.3, 18.6] 7.0 [3.4, 13.7] 7.0 [3.4, 13.7] 2.0 [0.6, 7.0] 3.0 [1.0, 8.5] 9.0 [4.8, 16.2] 5.0 [2.2, 11.2] 6.6 [3.3, 13.2]
Role 14.3 [7.9, 24.3] 12.9 [6.9, 22.7] 7.1 [3.1, 15.7] 20.6 [12.7, 31.6] 7.4 [3.2, 16.1] 11.8 [6.1, 21.5] 16.2 [9.3, 26.7] 19.1 [11.5, 30.0] 13.7 [7.6, 23.6]
Morphological Derivation 0.0 [0.0, 0.0] 2.2 [0.6, 7.5] 1.1 [0.2, 5.8] 4.3 [1.7, 10.5] 5.4 [2.3, 12.0] 1.1 [0.2, 5.8] 5.4 [2.3, 12.0] 6.5 [3.0, 13.4] 3.2 [1.3, 8.4]
Compound 6.0 [2.8, 12.5] 4.0 [1.6, 9.8] 3.0 [1.0, 8.5] 4.0 [1.6, 9.8] 7.0 [3.4, 13.7] 1.0 [0.2, 5.4] 8.0 [4.1, 15.0] 6.0 [2.8, 12.5] 4.9 [2.2, 10.9]
Discourse 5.7 [2.5, 12.8] 9.2 [4.7, 17.1] 3.4 [1.2, 9.7] 3.4 [1.2, 9.7] 2.3 [0.6, 8.0] 3.4 [1.2, 9.7] 5.7 [2.5, 12.8] 5.7 [2.5, 12.8] 4.9 [2.0, 11.5]
Average 8.8 [5.0, 15.2] 8.7 [4.9, 15.6] 7.8 [4.2, 14.6] 8.3 [4.6, 15.2] 7.1 [3.9, 13.7] 5.1 [2.7, 11.0] 8.4 [4.6, 15.4] 8.2 [4.5, 15.0] 7.8 [4.3, 14.5]
Table 15: Dialogue: Unrobustness (U, %) by model and modification
On Sentiment Analysis task, overall, PLMs were less robust to modifications compared to LLMs (see Table˜16), with the exception of Claude-3.5 which lacks robusteness to many tests that the other LLMs are stable on. Most prominently, all LLMs and PLMs are significantly unrobust to negation, and can be easily misguided by adding another appraisal marker to a sentence already bearing some sentiment.
Category Modification PLM LLM
BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 5.0 [2.2, 11.2] 5.0 [2.2, 11.2] 3.0 [1.0, 8.5] 0.0 [0.0, 0.0] 8.0 [4.1, 15.0] 1.0 [0.2, 5.4] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 3.5 [1.5, 8.5]
Geographical 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 3.0 [1.0, 8.5] 9.0 [4.8, 16.2] 8.0 [4.1, 15.0] 7.0 [3.4, 13.7] 4.0 [1.6, 9.8] 8.0 [4.1, 15.0] 6.0 [2.8, 12.4]
Length 3.0 [1.0, 8.5] 2.0 [0.6, 7.0] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 9.0 [4.8, 16.2] 2.0 [0.6, 7.0] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 3.0 [1.2, 8.3]
Orthographic Spelling 4.0 [1.6, 9.8] 4.0 [1.6, 9.8] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 6.0 [2.8, 12.5] 3.0 [1.0, 8.5] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 2.1 [0.9, 5.1]
Capitalization 0.0 [0.0, 0.0] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 1.0 [0.2, 5.5] 7.1 [3.5, 13.9] 3.0 [1.0, 8.5] 0.0 [0.0, 0.0] 3.0 [1.0, 8.5] 2.5 [1.0, 6.7]
Punctuation 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 0.0 [0.0, 0.0] 1.0 [0.2, 5.4] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 1.5 [0.4, 5.7]
Morphological Derivation 3.4 [1.2, 9.7] 5.7 [2.5, 12.8] 2.3 [0.6, 8.0] 2.3 [0.6, 8.0] 6.9 [3.2, 14.2] 2.3 [0.6, 8.0] 3.4 [1.2, 9.7] 4.6 [1.8, 11.2] 3.9 [1.5, 10.2]
Compound 3.2 [1.1, 8.9] 8.4 [4.3, 15.7] 3.2 [1.1, 8.9] 2.1 [0.6, 7.4] 5.3 [2.3, 11.7] 1.1 [0.2, 5.7] 3.2 [1.1, 8.9] 4.2 [1.6, 10.3] 3.8 [1.5, 9.7]
Syntactic Voice 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 9.0 [4.8, 16.2] 4.0 [1.6, 9.8] 10.0 [5.5, 17.4] 3.0 [1.0, 8.5] 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 6.8 [3.4, 13.3]
Grammar 4.5 [1.6, 12.5] 12.1 [6.3, 22.1] 3.0 [0.8, 10.4] 4.5 [1.6, 12.5] 7.6 [3.3, 16.5] 4.5 [1.6, 12.5] 4.5 [1.6, 12.5] 4.5 [1.6, 12.5] 5.7 [2.3, 14.0]
Conjunction 3.0 [1.0, 8.5] 6.0 [2.8, 12.5] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 3.0 [1.0, 8.5] 3.0 [1.0, 8.5] 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 2.1 [0.8, 6.1]
Semantic Concept 6.0 [2.8, 12.5] 6.0 [2.8, 12.5] 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 4.0 [1.6, 9.8] 2.0 [0.6, 7.0] 1.0 [0.2, 5.4] 5.0 [2.2, 11.2] 4.1 [1.7, 9.9]
Negation 22.9 [15.6, 32.3] 20.8 [13.9, 30.0] 25.0 [17.4, 34.5] 16.7 [10.5, 25.4] 17.7 [11.4, 26.5] 15.6 [9.7, 24.2] 16.7 [10.5, 25.4] 17.7 [11.4, 26.5] 19.1 [12.6, 28.1]
Pragmatic Discourse 3.0 [1.0, 8.5] 6.1 [2.8, 12.6] 4.0 [1.6, 9.9] 1.0 [0.2, 5.5] 12.1 [7.1, 20.0] 3.0 [1.0, 8.5] 2.0 [0.6, 7.1] 3.0 [1.0, 8.5] 4.3 [1.9, 10.1]
Sentiment 19.0 [12.5, 27.8] 16.0 [10.1, 24.4] 12.0 [7.0, 19.8] 14.0 [8.5, 22.1] 15.0 [9.3, 23.3] 16.0 [10.1, 24.4] 11.0 [6.3, 18.6] 10.0 [5.5, 17.4] 14.1 [8.7, 22.2]
Genre Casual 9.0 [4.8, 16.2] 9.0 [4.8, 16.2] 5.0 [2.2, 11.2] 3.0 [1.0, 8.5] 7.0 [3.4, 13.7] 6.0 [2.8, 12.5] 3.0 [1.0, 8.5] 2.0 [0.6, 7.0] 5.5 [2.6, 11.7]
Dialectal 9.8 [5.4, 17.1] 9.8 [5.4, 17.1] 7.8 [4.0, 14.7] 4.9 [2.1, 11.0] 6.9 [3.4, 13.5] 5.9 [2.7, 12.2] 2.9 [1.0, 8.3] 3.9 [1.5, 9.7] 6.5 [3.2, 12.9]
Average 6.5 [3.5, 12.7] 7.6 [4.0, 14.5] 5.3 [2.7, 10.8] 4.2 [2.0, 9.5] 8.0 [4.2, 15.1] 4.8 [2.3, 10.8] 3.6 [1.7, 8.5] 4.4 [2.1, 9.8] 5.6 [2.8, 11.5]
Table 16: Sentiment Analysis: Unrobustness (U, %) by model and modification
A.9 Use Case: GSM and IFEval (extended results)
We present the full results of GSM and IFEval tasks with CI intervals in Table˜17 and Table˜18.
Category Modification GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 4.0 [1.6, 9.8] 0.0 [0.0, 0.0] 1.0 [0.2, 5.4] 1.6 [0.5, 5.5]
Geographical 5.0 [2.2, 11.2] 5.0 [2.2, 11.2] 7.0 [3.4, 13.7] 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 4.0 [1.7, 9.7]
Length 4.0 [1.6, 9.8] 2.0 [0.6, 7.0] 2.0 [0.6, 7.0] 0.0 [0.0, 0.0] 2.0 [0.6, 7.0] 2.0 [0.6, 6.2]
Orthographic Spelling 1.0 [0.2, 5.4] 0.0 [0.0, 0.0] 2.0 [0.6, 7.0] 1.0 [0.2, 5.4] 0.0 [0.0, 0.0] 0.8 [0.2, 3.6]
Capitalization 3.0 [1.0, 8.5] 1.0 [0.2, 5.4] 5.0 [2.2, 11.2] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 2.2 [0.7, 7.2]
Punctuation 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 5.0 [2.2, 11.2] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 1.4 [0.5, 4.4]
Semantic Concept 1.0 [0.2, 5.4] 0.0 [0.0, 0.0] 5.0 [2.2, 11.2] 2.0 [0.6, 7.0] 1.0 [0.2, 5.4] 1.8 [0.6, 5.8]
Negation 15.0 [9.3, 23.3] 15.0 [9.3, 23.3] 18.0 [11.7, 26.7] 7.0 [3.4, 13.7] 9.0 [4.8, 16.2] 12.8 [7.7, 20.6]
Discourse Appraisal 0.0 [0.0, 0.0] 0.0 [0.0, 0.0] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 0.6 [0.1, 3.3]
Varieties Style 4.0 [1.6, 9.8] 3.0 [1.0, 8.5] 6.0 [2.8, 12.5] 4.0 [1.6, 9.8] 3.0 [1.0, 8.5] 4.0 [1.6, 9.8]
Dialect 2.0 [0.6, 7.0] 4.0 [1.6, 9.8] 6.0 [2.8, 12.5] 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 3.0 [1.1, 8.4]
Syntactic Conjunction 0.0 [0.0, 0.0] 1.0 [0.2, 5.4] 4.0 [1.6, 9.8] 1.0 [0.2, 5.4] 1.0 [0.2, 5.4] 1.4 [0.4, 5.2]
Voice 2.0 [0.6, 7.0] 2.0 [0.6, 7.0] 6.0 [2.8, 12.5] 1.0 [0.2, 5.4] 2.0 [0.6, 7.0] 2.6 [0.9, 7.8]
Average 2.9 [1.3, 7.1] 2.7 [1.2, 6.5] 5.5 [2.6, 11.6] 1.6 [0.5, 5.7] 2.0 [0.7, 6.6] 2.9 [1.3, 7.5]
Table 17: GSM: Unrobustness (U, %) by model and modification
Category Modification GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 11.1 [6.3, 18.8] 3.0 [1.0, 8.5] 11.1 [6.3, 18.8] 5.1 [2.2, 11.3] 12.1 [7.1, 20.0] 8.5 [4.6, 15.5]
Geographical 9.0 [4.8, 16.2] 12.0 [7.0, 19.8] 14.0 [8.5, 22.1] 9.0 [4.8, 16.2] 10.0 [5.5, 17.4] 10.8 [6.1, 18.4]
Length 11.0 [6.3, 18.6] 10.0 [5.5, 17.4] 17.0 [10.9, 25.5] 6.0 [2.8, 12.5] 16.0 [10.1, 24.4] 12.0 [7.1, 19.7]
Orthography Capitalization 8.1 [4.2, 15.1] 5.1 [2.2, 11.3] 10.1 [5.6, 17.6] 9.1 [4.9, 16.4] 14.1 [8.6, 22.3] 9.3 [5.1, 16.6]
Punctuation 7.1 [3.5, 13.9] 3.0 [1.0, 8.5] 13.1 [7.8, 21.2] 8.1 [4.2, 15.1] 12.1 [7.1, 20.0] 8.7 [4.7, 15.7]
Spelling 4.1 [1.6, 10.1] 4.1 [1.6, 10.1] 7.2 [3.5, 14.2] 6.2 [2.9, 12.8] 8.2 [4.2, 15.4] 6.0 [2.8, 12.5]
Syntax Conjunction 11.0 [6.3, 18.6] 8.0 [4.1, 15.0] 6.0 [2.8, 12.5] 8.0 [4.1, 15.0] 11.0 [6.3, 18.6] 8.8 [4.7, 15.9]
Voice 9.0 [4.8, 16.2] 9.0 [4.8, 16.2] 11.0 [6.3, 18.6] 6.0 [2.8, 12.5] 15.0 [9.3, 23.3] 10.0 [5.6, 17.4]
Semantics Concept 9.1 [4.9, 16.4] 7.1 [3.5, 13.9] 11.1 [6.3, 18.8] 7.1 [3.5, 13.9] 8.1 [4.2, 15.1] 8.5 [4.5, 15.6]
Negation 25.0 [17.5, 34.3] 22.0 [15.0, 31.1] 24.0 [16.7, 33.2] 23.0 [15.8, 32.2] 24.0 [16.7, 33.2] 23.6 [16.4, 32.8]
Discourse Appraisal 7.1 [3.5, 14.0] 4.1 [1.6, 10.0] 12.2 [7.1, 20.2] 7.1 [3.5, 14.0] 6.1 [2.8, 12.7] 7.3 [3.7, 14.2]
Varieties Style 18.0 [11.7, 26.7] 7.0 [3.4, 13.7] 12.0 [7.0, 19.8] 7.0 [3.4, 13.7] 11.0 [6.3, 18.6] 11.0 [6.4, 18.5]
Dialect 12.0 [7.0, 19.8] 13.0 [7.8, 21.0] 12.0 [7.0, 19.8] 11.0 [6.3, 18.6] 15.0 [9.3, 23.3] 12.6 [7.5, 20.5]
Average 10.9 [6.3, 18.4] 8.3 [4.5, 15.1] 12.4 [7.4, 20.2] 8.7 [4.7, 15.7] 12.5 [7.5, 20.4] 10.5 [6.1, 17.9]
Table 18: IFEval: Unrobustness (U, %) by model and modification
A.10 Scaling analysis (extended results)
We present the scaling analysis results with confidence intereval in Table˜19.
Category Modification Llama-8b Llama-70b Llama-405b Avg
Bias Temporal 8.0 [4.1, 15.0] 5.0 [2.2, 11.2] 4.0 [1.6, 9.8] 5.7 [2.6, 12.0]
Geographical 8.0 [4.1, 15.0] 3.0 [1.0, 8.5] 7.0 [3.4, 13.7] 6.0 [2.9, 12.4]
Length 8.0 [4.1, 15.0] 2.0 [0.6, 7.0] 2.0 [0.6, 7.0] 4.0 [1.7, 9.7]
Orthographic Spelling 10.0 [5.5, 17.4] 2.0 [0.6, 7.0] 2.0 [0.6, 7.0] 4.7 [2.2, 10.5]
Capitalization 11.0 [6.3, 18.6] 2.0 [0.6, 7.0] 5.0 [2.2, 11.2] 6.0 [3.0, 12.3]
Punctuation 5.0 [2.2, 11.2] 1.0 [0.2, 5.4] 5.0 [2.2, 11.2] 3.7 [1.5, 9.3]
Semantic Concept 12.0 [7.0, 19.8] 1.0 [0.2, 5.4] 5.0 [2.2, 11.2] 6.0 [3.1, 12.1]
Negation 22.0 [15.0, 31.1] 17.0 [10.9, 25.5] 18.0 [11.7, 26.7] 19.0 [12.5, 27.8]
Discourse Appraisal 7.0 [3.4, 13.7] 5.0 [2.2, 11.2] 1.0 [0.2, 5.4] 4.3 [1.9, 10.1]
Varieties Style 16.0 [10.1, 24.4] 6.0 [2.8, 12.5] 6.0 [2.8, 12.5] 9.3 [5.2, 16.5]
Dialect 10.0 [5.5, 17.4] 6.0 [2.8, 12.5] 6.0 [2.8, 12.5] 7.3 [3.7, 14.1]
Syntactic Conjunction 12.0 [7.0, 19.8] 2.0 [0.6, 7.0] 4.0 [1.6, 9.8] 6.0 [3.0, 12.2]
Voice 11.0 [6.3, 18.6] 2.0 [0.6, 7.0] 6.0 [2.8, 12.5] 6.3 [3.2, 12.7]
Average 10.8 [6.2, 18.2] 4.2 [1.9, 9.8] 5.5 [2.6, 11.6] 6.8 [3.6, 13.2]
Table 19: GSM Scaling: Unrobustness (U, %) by Llama model size and modification
A.11 Unrobustness results for each model averaged across tasks
Table˜20 reports the Unrobustness scores for each model averaged across all tasks. Note that IFEval and GSM tasks only contain LLMs results. Overall, PLMs are more unrobust compared to LLMs. Among LLMs, the newer reasoning models are more robust compared to the base versions. Modification-wise, Geographical Bias, Negation, Style, and Dialect are the most prominent ones in creating instability.
Category Modification BERT GPT-2 T5 GPT-4o Claude-3.5 Llama 3.1 GPT-5 DS R1 Avg
Bias Temporal 6.4 [3.1, 12.3] 4.5 [1.8, 10.0] 4.3 [1.9, 9.3] 4.7 [2.3, 9.5] 3.9 [1.5, 9.2] 4.3 [2.0, 9.5] 3.6 [1.6, 8.0] 6.1 [3.2, 11.9] 4.7 [2.2, 10.0]
Geographical 11.2 [7.2, 17.6] 13.6 [9.1, 20.5] 13.2 [9.1, 19.9] 11.4 [7.1, 18.3] 11.7 [7.3, 18.8] 12.0 [7.6, 18.9] 10.1 [6.3, 16.5] 11.8 [7.7, 18.5] 11.9 [7.7, 18.6]
Length 11.0 [6.4, 17.9] 8.8 [5.0, 15.2] 8.8 [5.1, 15.1] 8.3 [4.6, 14.8] 10.5 [6.1, 17.6] 7.6 [4.4, 13.5] 5.1 [2.5, 10.1] 8.6 [5.0, 15.0] 8.6 [4.9, 14.9]
Orthographic Spelling 5.0 [2.2, 11.2] 5.0 [2.2, 11.2] 2.0 [0.8, 4.9] 3.0 [1.4, 6.8] 3.0 [1.3, 7.0] 2.3 [0.7, 7.5] 0.7 [0.1, 3.6] 2.0 [0.9, 4.2] 2.9 [1.2, 7.0]
Capitalization 0.0 [0.0, 0.0] 4.6 [1.9, 10.8] 4.6 [1.9, 10.8] 2.4 [0.8, 7.6] 3.4 [1.4, 8.9] 3.4 [1.3, 9.0] 1.7 [0.6, 5.2] 1.7 [0.5, 6.5] 2.7 [1.1, 7.3]
Punctuation 2.5 [0.9, 7.6] 2.0 [0.6, 7.0] 4.0 [2.0, 7.5] 1.3 [0.4, 4.6] 2.0 [0.7, 5.7] 3.0 [1.1, 8.4] 2.3 [0.9, 7.3] 1.3 [0.3, 5.9] 2.3 [0.9, 6.8]
Semantic Concept 6.5 [3.1, 13.1] 5.5 [2.5, 11.8] 4.5 [1.9, 10.5] 4.3 [2.0, 10.1] 3.7 [1.7, 7.8] 3.0 [1.1, 8.4] 4.3 [2.1, 9.9] 4.0 [1.7, 9.7] 4.5 [2.0, 10.2]
Negation 25.9 [18.3, 35.4] 29.9 [21.9, 39.4] 28.0 [20.1, 37.5] 20.9 [14.2, 29.8] 20.9 [14.2, 29.8] 21.2 [14.4, 30.2] 16.2 [10.5, 24.5] 17.2 [11.2, 25.7] 22.5 [15.6, 31.5]
Discourse Appraisal 5.2 [2.5, 9.9] 5.3 [2.4, 10.0] 4.4 [2.0, 8.8] 4.6 [2.2, 8.6] 3.5 [1.5, 7.4] 6.3 [3.2, 11.8] 5.0 [2.5, 10.2] 7.0 [3.9, 12.8] 5.2 [2.6, 9.9]
Varieties Style 13.9 [8.3, 21.0] 15.6 [9.9, 22.8] 15.6 [10.1, 22.5] 11.2 [6.9, 17.5] 7.2 [3.8, 13.1] 9.6 [5.3, 16.2] 8.1 [4.4, 14.2] 9.2 [5.3, 15.6] 11.3 [6.8, 17.9]
Dialect 8.3 [4.4, 15.1] 26.9 [19.4, 35.9] 22.2 [15.4, 30.9] 12.8 [8.3, 20.2] 9.6 [5.4, 16.7] 11.6 [6.8, 19.1] 5.0 [2.5, 10.8] 9.9 [5.8, 17.0] 13.3 [8.5, 20.7]
Syntactic Conjunction 4.0 [1.6, 9.8] 5.0 [2.2, 11.2] 4.0 [1.8, 9.6] 2.3 [1.0, 6.0] 1.7 [0.5, 6.4] 3.0 [1.1, 8.4] 1.7 [0.6, 5.1] 1.7 [0.6, 5.1] 2.9 [1.2, 7.7]
Voice 9.0 [4.8, 16.2] 10.5 [5.9, 18.0] 8.0 [4.1, 14.9] 4.3 [1.9, 10.2] 4.7 [2.2, 10.5] 4.0 [1.6, 9.8] 5.0 [2.4, 10.9] 3.7 [1.5, 9.3] 6.1 [3.0, 12.5]
Average 8.4 [4.8, 14.4] 10.6 [6.5, 17.2] 9.5 [5.9, 15.5] 7.0 [4.1, 12.6] 6.6 [3.7, 12.2] 7.0 [3.9, 13.1] 5.3 [2.9, 10.5] 6.5 [3.7, 12.1] 7.6 [4.4, 13.5]
Table 20: All Tasks: Unrobustness (U, %) by model and modification (averaged across tasks)
A.12 AI usage disclosure
Claude-4-Sonnet (through Claude Code) was used to help debug issues in the code and help generating the tables in LaTex.